OpenStudy (anonymous):

one canned juice drink is 20% orange juice, another is 10% O.J. How many liters of each should be mixed together to get 10L that is 11% O.J....please help!!!!!

jimthompson5910 (jim_thompson5910):

let x = amount of 20% OJ y = amount of 10% OJ

jimthompson5910 (jim_thompson5910):

you need 10 L total, so x+y = 10 solve for y to get y = 10 - x

jimthompson5910 (jim_thompson5910):

if you have x liters of 20% OJ, you will have 0.2x liters of pure OJ (and no water) if you have y liters of 10% OJ, you will have 0.1y liters of pure OJ (and no water)

jimthompson5910 (jim_thompson5910):

combine the two and you'll have 0.2x+0.1y liters of pure OJ you want a mix that's 11% OJ and you want 10 L of this mix, so you want 0.11*10 = 1.1 liters of pure OJ

jimthompson5910 (jim_thompson5910):

This means that your second equation is 0.2x + 0.1y = 1.1

OpenStudy (anonymous):

x+y=10 0.20x+0.10y=(10)(0.11) 20x+10x=110

jimthompson5910 (jim_thompson5910):

if you multiply everything in 0.2x + 0.1y = 1.1 by 10, you will get 0.2x + 0.1y = 1.1 10*0.2x + 10*0.1y = 10*1.1 2x + y = 11

jimthompson5910 (jim_thompson5910):

Now you plug in y = 10 - x 2x + y = 11 2x + 10-x = 11 and you would solve for x

OpenStudy (anonymous):

x+y=11 -20(x+y)=-20(11) -20x-20y=-220

jimthompson5910 (jim_thompson5910):

so what do you get when you solve 2x + 10-x = 11

OpenStudy (anonymous):

-8=11so -8/11?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

2x + 10-x = 11 x + 10 = 11 x = ??

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

x+y = 10 1+y = 10 y = 10-1 y = 9

jimthompson5910 (jim_thompson5910):

So you need 1 liter of the 20% OJ 9 liters of the 10% OJ